OpenStudy (anonymous):

WILL FAN AND MEDAL!! Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (igreen):

\(y \le x - 5\) \(y \ge -x - 4\) We can plug the points in and see which one fits both of these inequalities.

OpenStudy (anonymous):

\[-5 \le 2 - 5?\]

OpenStudy (igreen):

No, you got the 'y' and 'x' values mixed up

OpenStudy (anonymous):

oops put in x for y

OpenStudy (igreen):

Let's try (5, -2).

OpenStudy (igreen):

\(y \le x - 5\) \(-2 \le 5 - 5\) Simplify

OpenStudy (anonymous):

I'm sorry I'm really confused on this

OpenStudy (igreen):

Just subtract 5 - 5 and see if it's true.

OpenStudy (igreen):

5 - 5 = ?

OpenStudy (anonymous):

0

OpenStudy (igreen):

Yes, so we have: \(-2 \le 0\) Is -2 less than or equal to 0?

OpenStudy (anonymous):

less than

OpenStudy (igreen):

Yes, so this is true.

OpenStudy (igreen):

Now let's plug in (5, -2) into the 2nd inequality to see if that fits also. \(y \ge -x - 4\) \(-2 \ge -5 - 4\) Again, subtract -5 - 4

OpenStudy (anonymous):

-9

OpenStudy (igreen):

Yes, now we have: \(-2 \ge -9\) Is -2 greater than -9?

OpenStudy (anonymous):

yes :)

OpenStudy (igreen):

Yes, so (5, -2) fits both of the inequalities, so it is our answer.

OpenStudy (anonymous):

Thank you so much I'm understanding this

OpenStudy (igreen):

Good to know :)

OpenStudy (anonymous):

I have a's in all subjects except for math so thanks you've been a lifesaver <3

OpenStudy (igreen):

No problem, I always have A's in all my courses, but I hate all of them except for Math :P